OpenStudy (anonymous):

What is the value of x in the equation below? 2 + 3(5 − 4x) = 17 + 4(4x − 1)

OpenStudy (anonymous):

x=1/7

OpenStudy (anonymous):

how

OpenStudy (raden):

you have to grouped them by combine the similar terms. 2 + 3(5 − 4x) = 17 + 4(4x − 1) 2 + 15 − 12x = 17 + 16x − 4 17 - 12x = 13 + 16x -12x - 16x = 13 - 17 -28x = -4 x = -4/-28 x = 1/7

OpenStudy (anonymous):

ok thank you soooo much

OpenStudy (raden):

you're welcome